# Multi-Conditional Ranking with Large Language Models

### v1

@@ Abstract
Utilizing large language models (LLMs) to rank a set of items has become a common approach in recommendation and retrieval systems. Typically, these systems focus on ordering a substantial number of documents in a monotonic order based on a given query. However, real-world scenarios often present a different challenge: ranking a comparatively smaller set of items, but according to a variety of diverse and occasionally conflicting conditions. In this paper, we define and explore the task of multi-conditional ranking by introducing MCRank, a benchmark tailored for assessing multi-conditional ranking across various item types and conditions. Our analysis of LLMs using MCRank indicates a significant decrease in performance as the number and complexity of items and conditions grow. To overcome this limitation, we propose a novel decomposed reasoning method, consisting of EXtracting and Sorting the conditions, and then Iterativly Ranking the items (EXSIR). Our extensive experiments show that this decomposed reasoning method enhances LLMs’ performance significantly, achieving up to a 12% improvement over existing LLMs. We also provide a detailed analysis of LLMs performance across various condition categories, and examine the effectiveness of decomposition step. Furthermore, we compare our method with existing approaches such as Chain-of-Thought and an encoder-type ranking model, demonstrating the superiority of our approach and complexity of MCR task. We released our dataset and code [^1] https://github.com/megagonlabs/MCR.
@@ 1 Introduction
The rapid advancement of autoregressive Large Language Models (LLMs) has significantly enhanced our ability to understand and solve NLP related tasks (Chowdhery et al., 2022; Touvron et al., 2023; OpenAI, 2023; Team et al., 2023). Among these tasks, document ranking plays a crucial role in recommendation and retrieval systems(Wu et al., 2023; Zhu et al., 2023). While there has been a considerable advancement in ranking extensive document collections given a query (Khattab & Zaharia, 2020; Zhuang et al., 2023b; Qin et al., 2023), the nuanced task of ranking a smaller set of items based on multiple conditions—a critical requirement in numerous real-world applications—has not been addressed in prior research.
Ranking a set of items according to multiple conditions has vast implications across various fields and applications. In recommendation systems, for instance, once the top candidates are shortlisted, the user experience can be significantly enhanced by offering the capability to re-rank these candidates based on specific conditions, such as genres and categories. In the realm of education, this task can be applied to the ranking of questions, enabling educators to prioritize and arrange questions effectively according to different criteria, such as subject matter. Moreover, in the competitive environment of job markets, multi-conditional ranking is invaluable for aligning resumes with job postings, while prioritizing various factors like certain skills and experience level.
@@
In this paper, we define and explore the task of multi-conditional ranking (MCR) by developing MCRank, a comprehensive benchmark consisting various item types and conditions for assessing MCR task. In addition, we also propose a novel decomposed reasoning based method, EXSIR, that beats strong baselines (including CoT) by up to 12%. The new benchmark, MCRank, spans a various category of conditions, including positional, locational, temporal, trait-based, and reasoning types. We have designed MCRank to address scenarios involving one to three conditions and to assess sets of 3, 5, or 7 items. The benchmark distinguishes between two types of items: token-level items, which consist of only a few tokens, and paragraph-level items, which contain up to 150 tokens. An example of MCRank involving two conditions and three token-level items is presented in Figure 1.
@@
Our initial investigations into the performance of existing LLMs on MCRank revealed a notable decline in accuracy as the number of items and conditions increased. Specifically, we observe that the accuracy of investigated LLMs , .i.e., GPT-4 (OpenAI, 2023), ChatGPT(Kocoń et al., 2023) (both turbo versions), and Mistral (Jiang et al., 2023) in correctly ranking items fell dramatically as the task scaled to three conditions and seven items, with accuracy approaching nearly 0%. To address the shortcomings of existing LLMs in MCR task, we introduce a novel method based on decomposed reasoning. Rather than directly prompting LLMs to rank items based on the given conditions, our approach begins with extracting and sorting the conditions based on their priority. Subsequently, we iteratively apply these sorted conditions to the item list. An illustration of our approach is provided in Figure 1. Applying our method to MCRank, we observed a notable improvement, with up to a 12% increase in the LLMs’ ranking accuracy.
@@
Observing the impact of our approach in improving LLMs performance on MCRank, we conduct an in-depth analysis of the models’ performance, dissecting the results based on the types of items and conditions involved. Additionally, we examine the accuracy of the decomposition step within the evaluated LLMs shedding light on observed behavior of LLMs on MCRank. To delve deeper into the significance of the decomposition process, we incorporate a zero-shot chain-of-thought (Wei et al., 2022) approach, further underscoring the importance of segmenting the MCR task into multiple steps to achieve improved outcomes. Finally, we employ ColBERT (Khattab & Zaharia, 2020), a model renowned for its efficacy in document ranking (Nguyen et al., 2016; Dietz et al., 2017), to represent encoder-type rankers. Our comparison illustrates that, despite their success in existing ranking tasks, encoder-type rankers exhibit considerably inferior performance on MCRank when compared to autoregressive LLMs.
@@
Positional Item “[one of the items]” should be the last from left
Locational Items that are in Africa should appear at the beginning
Temporal Sort items based on their deadline from the first to the last
Trait-Based Sort the items based on their size from the smallest to the largest
Reason-Based Items that has the largest yards of touchdown should appear at the beginning
Table 1: Example of conditions based on different types. After extracting items, we determine their golden ranking based on their corresponding labels. The conditions may specify a complete order for all items or only provide a partial ordering instructions for the placement of certain items.
2 Multi-Conditional Ranking
The task of multi-conditional ranking is designed to shift the focus from traditional ranking tasks, which typically involve ordering a large set of items based on a single query. Instead, this task concentrates on sorting a smaller, pre-selected set of items according to multiple conditions. These conditions may not only conflict with one another but also carry varying levels of priority, adding layers of complexity to the ranking task. Moreover, each condition may specify a complete order for all items or only provide a partial ordering instructions for the placement of certain items. The primary objective is to closely replicate a scenario in which a user provides a string containing several conditions along with their respective priorities, which then guides the ranking of an already shortlisted group of items. The complexity of this task lies in its need to balance various conditions, understand their relative importance, and apply them effectively to produce a nuanced and contextually relevant ordering of items.
@@ 2.1 MCRank Benchmark
To develop a benchmark for assessing the ability of large language models (LLMs) to tackle the multi-conditional ranking (MCR) task—where the goal is to rank a small set of items based on a string of unsorted conditions—we must first compile a collection of items, each tagged with a gold label that denotes a specific category or a value for a particular feature. These labels serve as the foundation for generating the correct ranking order under any given set of conditions. To structure our benchmark, we classify the conditions into five distinct types and distinguish items based on two categories: (1) token-level, which includes items comprising only a few tokens, and (2) paragraph-level, which encompasses items containing up to 150 tokens. We then collect items and their corresponding labels for each category. The conditions are divided into the categories below. The detailed list of samples for each category and the datasets used is provided in the Appendix. We also provide an example for each condition category in Table 1.
@@
Locational: Locational conditions are defined as conditions that require the placement of items in the ranking based on their geographical attributes. For the token-level category, we compile items and their respective locational labels by extracting popular entities and their objects from the location predicate found in the T-REx benchmark (Elsahar et al., 2018). Additionally, for the paragraph-level category, to encompass a broader range of items, we combine prompts for birth place, death place, country of citizenship, headquarter location, and location predicates from the T-REx benchmark along with job descriptions from Dice  [^2] We extracted the data from https://www.kaggle.com/datasets/PromptCloudHQ/us-technology-jobs-on-dicecom that contain locational labels.
@@
Reason-Based: We define reason-based conditions as those that necessitate logical/mathematical reasoning to determine the correct positioning of items, such as deducing the category of an item or performing mathematical operations on provided values of a certain attribute in each item. For the token-level category, we collect items and their categories from the auto-categorization task in Big-Bench (Srivastava et al., 2022). In the paragraph-level category, we sourced items from DROP’s (Dua et al., 2019) paragraphs featuring ”How Many” questions that require mathematical reasoning.
@@
To develop the MCRank benchmark, we assemble a collection of datasets, each corresponding to one of the two item categories and featuring samples with 1, 2, or 3 conditions and sets of 3, 5, or 7 items, culminating in 18 distinct scenarios. We curate the dataset for each scenario through several steps: Initially, for each condition type, we compile data and their labels to create 200 samples. Each sample includes a condition from that category, a randomly arranged set of items, and the correct item ranking based on the labels. For the positional type conditions, we utilize items from the auto-categorization task in Big-Bench for the token-level and Amazon reviews for the paragraph-level category.
After assembling 200 samples for each type of condition across all scenarios, we introduce additional conditions for scenarios requiring more than one condition to mimic a realistic setting where users specify various conditions. We randomly add either a condition to sort items based on character counts or a positional condition. In scenarios with three conditions, we incorporate both. We assign a “low priority” to the character count condition, a “medium priority” to each category type condition, and a “high priority” to the extra positional condition, aiming to reflect the varying priorities a user might have in real-world contexts for these conditions. Subsequently, we randomize the conditions order and combine the samples from each condition type to form a dataset for each scenario. To maintain clarity, we eliminate samples where multiple items share the same character counts. The detailed statistics of MCRank are presented in Section 2. As indicated, we curated approximately 930 samples on average for each scenario.
1 Condition 2 Conditions 3 Conditions
Token-level 916.7 860.0 797.7
Paragraph-level 1000 1000 1000
Table 2: Average number of samples in MCRank benchmark per number of conditions.
2.2 Extracting, Sorting, and Iterativly Ranking (EXSIR)
@@
3 Experimental Details
Models
We evaluate both commercial LLMs, GPT-4 (OpenAI, 2023), ChatGPT(Kocoń et al., 2023) (both turbo versions), as well as the open-source LLM, Mistral (Jiang et al., 2023) (Mistral-7B-Instruct-v0.2) on MCRank. To address the MCR task, in the base setting, we input the string of conditions along with the list of items into the prompt, instructing the LLMs to organize the list according to the conditions. For paragraph-level items, we assign a unique label, ”Item-K,” to each item. The task for the model is then to rank the items but to output the sequence of sorted labels—Item-K—instead of the items themselves. The details of all prompts utilized in this study are provided in the Appendix.
Evaluation Metric
Given that the MCR task defined in this paper represents a broader and more complex variation of previously defined ranking tasks—where, unlike those tasks, the significance or relevance of items in the gold ranking doesn’t necessarily diminish in a linear order—traditional ranking metrics like MRR or nDCG (Zangerle & Bauer, 2022) are not suitable for our context. Consequently, we evaluate model performance on the MCR task using exact match accuracy, where a completely correct ranking earns a score of 1, and an incorrect one receives 0. Additionally, we employ an averaged accuracy metric, calculating the mean number of items correctly positioned in each sample to provide a nuanced view of the models’ ranking abilities.
4 Experiments
In this section, we explore the effectiveness of LLMs and the influence of the EXISR method on the MCR task using the MCRank benchmark. Our analysis begins with an evaluation of the models’ performance on MCRank. Subsequently, we delve into a detailed breakdown of performance across different categories to assess each model’s capabilities. To understand the impact of the decomposition process on EXSIR’s functionality, we examine the accuracy of the decomposition step for each model. Lastly, to underscore the importance of decomposed reasoning through multiple steps, we compare our method’s performance against zero-shot Chain-of-Thought prompting and an encoder-based ranking model.
@@ 4.2 Per-Category Breakdown
A detailed breakdown of LLMs’ performance across different condition categories on the MCRank benchmark is available in the Appendix. Notably, performance varies across condition categories for each setting. In token-level scenarios, models excel in reason-based conditions, whereas, in paragraph-level settings, they perform better in locational conditions but exhibit comparable results in trait-based conditions. This variation is attributed to the increased complexity of reason-based conditions in paragraph settings and the explicit provision of label information in trait-based and locational conditions, simplifying these tasks. However, all models struggle with the positional conditions, especially with conditions like “item [x] should be the last from the right” and “the last item in the sorted list should appear in the first place.” This struggle is likely due to the conflict between these conditions and the autoregressive nature of the models, which necessitates a complete understanding of the final ranking before even generating the initial item.

### v3

@@ Abstract
Utilizing large language models (LLMs) to rank a set of items has become a common approach in recommendation and retrieval systems. Typically, these systems focus on ordering a substantial number of documents in a monotonic order based on a given query. However, real-world scenarios often present a different challenge: ranking a comparatively smaller set of items, but according to a variety of diverse and occasionally conflicting conditions. In this paper, we define and explore the task of multi-conditional ranking by introducing MCRank, a benchmark tailored for assessing multi-conditional ranking across various item types and conditions. Our analysis of LLMs using MCRank indicates a significant decrease in performance as the number and complexity of items and conditions grow. To overcome this limitation, we propose a novel decomposed reasoning method, consisting of EXtracting and Sorting the conditions, and then Iteratively Ranking the items (EXSIR). Our extensive experiments show that this decomposed reasoning method enhances LLMs’ performance significantly, achieving up to a 14.4% improvement over existing LLMs. We also provide a detailed analysis of LLMs performance across various condition categories, and examine the effectiveness of decomposition step. Furthermore, we compare our method with existing approaches such as Chain-of-Thought and existing ranking models, demonstrating the superiority of our approach and complexity of MCR task. We released our dataset and code [^1] https://github.com/megagonlabs/MCR.
Multi-Conditional Ranking with Large Language Models
Pouya Pezeshkpour Megagon Labs pouya@megagon.ai                        Estevam Hruschka Megagon Labs estevam@megagon.ai
@@ 1 Introduction
The rapid advancement of autoregressive Large Language Models (LLMs) has significantly enhanced our ability to understand and solve NLP related tasks (Chowdhery et al., 2022; Touvron et al., 2023; OpenAI, 2023; Team et al., 2023). Among these tasks, document ranking plays a crucial role in recommendation and retrieval systems (Wu et al., 2023; Zhu et al., 2023). While there has been a considerable advancement in ranking extensive document collections given a query (Khattab and Zaharia, 2020; Zhuang et al., 2023b; Qin et al., 2023), the nuanced task of ranking a smaller set of items based on multiple conditions—a critical requirement in numerous real-world applications—has not been addressed in prior research.
Ranking a set of items according to multiple conditions has vast implications across various fields and applications. In recommendation systems, for instance, once the top candidates are shortlisted, the user experience can be significantly enhanced by offering the capability to re-rank these candidates based on specific conditions, such as genres and categories. In the realm of education, this task can be applied to the ranking of questions, enabling educators to prioritize and arrange questions effectively according to different criteria, such as subject matter. Moreover, in competitive job markets, multi-conditional ranking is crucial for matching resumes to job postings, prioritizing factors like skills and experience.
@@
In this paper, we define and explore the task of multi-conditional ranking (MCR) by developing MCRank, a comprehensive benchmark consisting various item types and conditions for assessing MCR task. In addition, we also propose a novel decomposed reasoning based method, EXSIR, that beats strong baselines (including CoT) by up to 14.4%. The new benchmark, MCRank, spans a various category of conditions, including positional, locational, temporal, trait-based, and reasoning types. We have designed MCRank to address scenarios involving one to three conditions and to assess sets of 3, 5, or 7 items. The benchmark distinguishes between two types of items: token-level items, which consist of only a few tokens, and paragraph-level items, which contain up to 150 tokens. An example of MCRank involving two conditions and three token-level items is presented in Figure 1.
@@
Our initial investigations into the performance of existing LLMs on MCRank revealed a notable decline in accuracy as the number of items and conditions increased. Specifically, we observe that the accuracy of investigated LLMs , i.e., OpenAI o1-mini, GPT-4 (OpenAI, 2023), ChatGPT(Kocoń et al., 2023) (both turbo versions), Llama3.1-70B (Touvron et al., 2023) and Mistral (7B) (Jiang et al., 2023) in correctly ranking items fell dramatically as the task scaled to three conditions and seven items, with accuracy approaching nearly 0%. To address the shortcomings of existing LLMs in MCR task, we introduce a novel method based on decomposed reasoning. Rather than directly prompting LLMs to rank items based on the given conditions, our approach begins with extracting and sorting the conditions based on their priority. Subsequently, we iteratively apply these sorted conditions to the item list. An illustration of our approach is provided in Figure 1. Applying our method to MCRank, we observed a notable improvement, with up to a 14.4% increase in the LLMs’ ranking accuracy.
@@
Observing the impact of our approach in improving LLMs performance on MCRank, we conduct an in-depth analysis of the models’ performance, dissecting the results based on the types of items and conditions involved. Additionally, we examine the accuracy of the decomposition step within the evaluated LLMs shedding light on observed behavior of LLMs on MCRank. To delve deeper into the significance of the decomposition process, we incorporate a zero-shot chain-of-thought (Wei et al., 2022) approach, further underscoring the importance of segmenting the MCR task into multiple steps to achieve improved outcomes. Finally, we employ SFR (Meng et al., 2024), ColBERT (Khattab and Zaharia, 2020) and RankGPT (based on GPT-4) (Sun et al., 2023), rankers renowned for their performance in document ranking (Nguyen et al., 2016; Dietz et al., 2017), to represent existing rankers. Our comparison illustrates that, despite their success in existing ranking tasks, they exhibit considerably inferior performance on MCRank.
2 Multi-Conditional Ranking
The task of multi-conditional ranking is designed to shift the focus from traditional ranking tasks, which typically involve ordering a large set of items based on a single query. Instead, this task concentrates on sorting a smaller, pre-selected set of items according to multiple conditions. These conditions may not only conflict with one another but also carry varying levels of priority, adding layers of complexity to the ranking task. Moreover, each condition may specify a complete order for all items or only provide a partial ordering instructions for the placement of certain items. The primary goal is to simulate a scenario where a user provides a string containing several conditions along with their respective priorities, to guide the ranking of a already shortlisted group of items. This task’s complexity lies in balancing conditions, understanding their relative importance, and applying them to create a contextually relevant items order.
@@ 2 Multi-Conditional Ranking
Position Item “[X]” should be the last from left
Location Items that are in Africa should appear at the
beginning
Temporal Sort items based on their deadline from the
first to the last
Trait Sort the items based on their size from the
smallest to the largest
Reason Items that has the largest yards of touchdown
should appear at the beginning
Table 1: Example of conditions based on different types. After extracting items, we determine their golden ranking based on their corresponding labels. The conditions may specify a complete or partial order for item placement.
@@ 2.1 MCRank Benchmark
To develop a benchmark for assessing the ability of LLMs to tackle the multi-conditional ranking (MCR) task—where the goal is to rank a small set of items based on a string of unsorted conditions—we must first compile a collection of items, each tagged with a gold label that denotes a specific category or a value for a particular feature. These labels serve as the foundation for generating the correct ranking order under any given set of conditions. We structure the benchmark by classifying conditions into five types and categorizing items as either: (1) token-level, which includes items comprising only a few tokens, and (2) paragraph-level, which encompasses items containing up to 150 tokens. We then collect items and their corresponding labels for each category. The conditions are divided into the categories below. We aimed to be as comprehensive as possible in choosing the categories, drawing insights from previous works on various tasks such as question answering (Dua et al., 2019), relation extraction (Pawar et al., 2017), text classification (Pang et al., 2008), and retrieval systems (Zhao et al., 2024). Our motivation for each category was to capture a broad range of ranking needs across various subfields. The detailed list of samples for each category and the datasets used is provided in the Appendix. We also provide an example for each condition category in Table 1.
@@
Locational: Locational conditions are defined as conditions that require the placement of items in the ranking based on their geographical attributes. For the token-level category, we compile items and their respective locational labels by extracting popular entities and their objects from the location predicate found in the T-REx benchmark (Elsahar et al., 2018). Additionally, for the paragraph-level category, we combine prompts for birth place, death place, country of citizenship, headquarter location, and location predicates from the T-REx benchmark, along with job descriptions from Dice [^2] We extracted the data from https://www.kaggle.com/datasets/PromptCloudHQ/us-technology-jobs-on-dicecom containing locational labels.
@@
Reason-Based: We define reason-based conditions as those that necessitate logical/mathematical reasoning to determine the correct positioning of items, such as deducing the category of an item or performing mathematical operations on values of a certain attribute in each item. For the token-level category, we collect items and their categories from the auto-categorization task in Big-Bench (Srivastava et al., 2022). In the paragraph-level category, we sourced items from DROP’s (Dua et al., 2019) paragraphs featuring "How Many" questions that require mathematical reasoning.
@@
To develop the MCRank benchmark, we assemble a collection of datasets, each corresponding to one of the two item categories and featuring samples with 1, 2, or 3 conditions and sets of 3, 5, or 7 items, culminating in 18 distinct scenarios. We curate the dataset for each scenario through several steps: Initially, for each condition type, we compile data and their labels to create 200 samples. Each sample includes a condition from that category, a randomly arranged set of items, and the correct item ranking based on the labels. For the positional type conditions, we utilize items from the auto-categorization task in Big-Bench for the token-level and Amazon reviews for the paragraph-level category. After assembling 200 samples for each type of condition across all scenarios, we introduce additional conditions for scenarios requiring more than one condition to mimic a realistic setting where users specify various conditions. We randomly add either a condition to sort items based on character counts or a positional condition. In scenarios with three conditions, we incorporate both.
As in Boutilier (2013), we consider the users’ priority as extra input, and assume they are explicitly provided by the users. As described in Schnabel et al. (2020); Boutilier (2013), the aim is to discover enough about the user utility function to recommend a good recommendation. Thus, we further assign a “low priority” to the character count condition, a “medium priority” to each category type condition, and a “high priority” to the extra positional condition. The main usage of assigning priorities is to handle contradictory conditions. Moreover, our goal in choosing this specific assignment of conditions and their priorities is to capture the varying priorities a user might have in real-world contexts Stray et al. (2024). Specifically, we use character count to represent low-priority and easy conditions that a user might inquire about. A medium-priority condition is covered by samples from different categories of conditions, while a high-priority condition represents a scenario in which a user asks to place the hardest/easiest, highest/lowest quality, or most/least qualified item either first (as most important) or last (to be ignored). We believe that this specific form of assignment, along with the randomness/diversity in selecting various conditions, can encompass a wide range of potential conditions posed by users in real-world scenarios.
To further validate our assignment of priorities, let’s see a concrete application scenario for MCR in recommendation systems (more scenarios in the Appendix): Imagine a user interacting with a movie recommender system. After receiving a list of potential matches, the user may not want to read through the plot of each movie. Instead, based on their current mood, they may wish to rank the movies according to various conditions: (1) Movies should be sorted by their length, with low priority, as the user generally prefers shorter films but doesn’t mind longer ones either; (2) The movies should be sorted by their IMDB score, with medium priority; and (3) The highest-rated movie should be placed at the end of the list, with high priority, because the user has already watched it.
Upon assigning the priorities, we randomize the conditions order, adding another layer of complexity to make the task more realistic, and combine the samples from each condition type to form a dataset for each scenario. To maintain clarity, we eliminate samples where multiple items share the same character counts. We curated approximately 1000 samples for each scenario. We provide MCRank’s statistics, more details, and a step by step illustration of MCRank creation in the Appendix.
2.2 Extracting, Sorting, and Iteratively Ranking (EXSIR)
@@
3 Experimental Details
Models: We evaluate both commercial LLMs, OpenAI o1-mini, GPT-4 (OpenAI, 2023), ChatGPT(Kocoń et al., 2023) (both turbo versions), as well as open-source LLMs, Llama3.1-70B (instruct) (Touvron et al., 2023) and Mistral (Mistral-7B-Instruct-v0.2) (Jiang et al., 2023) on MCRank. To address the MCR task, in the base setting, we input the string of conditions along with the list of items into the prompt, instructing the LLMs to organize the list according to the conditions. For paragraph-level items, we assign a unique label, "Item-K," to each item. The task for the model is then to rank the items but to output the sequence of sorted labels—Item-K—instead of the items themselves. The details of all prompts utilized in this study are provided in the Appendix. We designed the prompts to mirror previous ranking works, while avoiding framing the task as a standard document monotonic-ranking task.
Evaluation Metric: Given that the MCR task defined in this paper represents a broader and more complex variation of previously defined ranking tasks—where, unlike those tasks, the significance or relevance of items in the gold ranking doesn’t necessarily diminish in a linear order—traditional ranking metrics like MRR or nDCG (Zangerle and Bauer, 2022) are not suitable for our context. Consequently, we evaluate model performance on the MCR task using exact match accuracy, scoring 1 for a fully correct ranking and 0 for an incorrect one. Additionally, we employ an averaged accuracy metric, measuring the mean number of correctly ranked items per sample for a more nuanced assessment of models.
4 Experiments
We explore the effectiveness of LLMs and the impact of EXSIR on the MCR task using the MCRank benchmark, starting with an evaluation of models’ performances. Subsequently, we provide per-category breakdown of performances to assess the impact if each condition category. We assess the accuracy of the decomposition step for each model to gauge its impact on EXSIR’s functionality. Lastly, to underscore the importance of decomposed reasoning through multiple steps, we compare our method’s performance against zero-shot CoT prompting and existing ranking models.
@@ 4.2 Per-Category Breakdown
A detailed breakdown of LLMs’ performance across different condition categories on the MCRank benchmark is available in the Appendix. Notably, performance varies across condition categories for each setting. In token-level scenarios, models excel in reason-based conditions, whereas, in paragraph-level settings, they perform better in locational conditions but exhibit comparable results in trait-based conditions. This variation is due to the increased complexity of reason-based conditions in paragraph settings and the explicit provision of label information in trait-based and locational conditions, simplifying these tasks. However, all models struggle with the positional conditions, especially with conditions like “item [X] should be the last from the right” and “the last item in the sorted list should appear in the first place.” This struggle is likely due to the conflict between the conditions and the models’ autoregressive nature, which necessitates a full understanding of the final ranking before even generating the first item. We also provide the accuracy of LLMs in satisfying high priority conditions in Appendix.